OpenStudy (anonymous):

Which equation is related to this equation? w ÷ 9 = 14

OpenStudy (anonymous):

A. w = 14 ÷ 9 B. w = 14 • 9 C. w = 14 – 9 D. w = 14 + 9

OpenStudy (jojokiw3):

9 x 14 = w. Btw I can't read that what you just posted.

OpenStudy (anonymous):

Its fine lol it was right thank you